# Spica: Retrieving Scenarios for Pluralistic In-Context Alignment

### v1

@@ 3.2.3 Learning Metric Weights
As our new metrics don’t replace relevance but only add group-level information, we need to balance potentially competing objectives. In Spica, we do this by taking a linear weighted combination of the introduced metrics and similarity (distance), in the form of: $\bar{d}(x,x^{\prime})=w_{d}\cdot d(x,x^{\prime})+w_{s}\cdot g_{\text{stability}}(x^{\prime})+w_{c}\cdot g_{\text{contrast}}(x^{\prime})+c$.
As the optimal weighting of metrics likely to vary across groups, we need to empirically find these weights. As discussed in Section 3.1, the retrieval desirability of $x^{\prime}$ given input $x$ can be assessed as an expected loss $\mathbb{E}[L(r(x),r(x^{\prime}))]$, we can compute the loss and minimize using a simple linear regression $\bar{d}(x,x^{\prime})=\mathbb{E}[L(r(x),r(x^{\prime}))]$. Here we show the case where only the best ($k=1$) is selected, though a similar approach could be applied for selecting a larger set of ${x^{\prime}}$.
@@ 3.3 In-Context Learning Prompts for Retrieved Scenarios
Because retrieved scenarios contain preference distributions across multiple responses (or strategies), different setups for integrating scenarios as demonstrations are likely to produce different model outputs. ICL prompt designs have been extensively studied by prior works, so in this work we primarily explore new configurations enabled by our construction of scenarios in the scenario bank.
Specifically, we focus on two factors: (1) the way preference distributions are incorporated into the prompt (The highest Positive preference or Contrasting pairs of highest and lowest preferences), and (2) the type of response feedback used (Instructions or example Response text). This results in 4 combinations of prompt setups: P-I, C-I, P-R, and C-R. Through human evaluations, we assess these prompting setups, with specific details of how we implement these prompts documented in Appendix A.6.2.
@@ 4 Experiments and Results
To evaluate Spica, we set up a small-scale proof-of-concept pluralistic alignment task involving 4 demographically constructed groups, and assess how well a Spica workflow is able to align model outputs to preferences of each group compared to a baseline approach that only considers semantic similarity. We conduct this evaluation over two language models and explore variations of the prompt used to incorporate scenarios. We assess the end-to-end alignment outputs and also examine the impact of pluralistic versus collective alignment.
@@ 4.1 Dataset and Scenario Bank Construction
For our evaluation alignment task, we constructed a set of queries (which define the topics to provide alignment on) by drawing from an existing set of challenging alignment situations based on prompts observed in conversations on the PRISM dataset Kirk et al. (2024). PRISM engaged human participants to interact with LLMs by naturally starting conversations with 3 types of guidance meant to invoke conversations around more challenging and complex topics: “unguided”, “values guided”, or “controversy guided”. We observed that of the 3 types of guidance, unguided conversations primarily resulted in simple informational requests which are not particularly controversial in the context of pluralistic alignment, so we opted to drop conversations of this type. Among the remaining conversations, we randomly selected a subset, split into 3 slices: retrieval (train, $n=360$), weight optimization and prompt strategy evaluation (dev, $n=150$), and evaluation hold-out (test, $n=75$).
For the responses used in our scenario bank, we did not use responses from PRISM, instead we follow Section 3.1 and construct responses for each aligned model ourselves based on a set of strategies covering common behaviors Appendix A.6.1). To capture the stochastic nature of model outputs, we generate 3 responses for each strategy.
@@ 4.2 Models and Similarity Metric
For our experiments, we tested the quality of retrieval-based ICL alignment using one open-source (llama3-8b) and one closed source model (gpt-4o-2024-05-13) as the base model. llama3-8b inference was conducted using a locally hosted instance of Ollama  [^2] https://ollama.com/. With both models, we applied the same prompts to generate responses attached to scenario bank queries and to conduct apply in-context alignment (Appendix A.6.2).
As our goal is to evaluate the additional metrics we introduced, we held the semantic similarity measurements constant across all models and conditions, using values derived by computing the cosine similarity between embeddings generated by text-embedding-3-large.
@@ 4.3 Pluralistic Groups and Human Annotation Setup
For our pluralistic alignment task, we define four groups in the form of demographic slices drawn from the US populations: partisan political affiliation (“Republican” or “Democrat”), and self-reported regular participation in religious activities (“yes” or “no”). Our choice of these specific features is based on these factors being salient for different opinions around AI Zhang and Dafoe (2019) along with practical considerations around available demographics recruitable on our crowd platform Prolific.
Annotators in each group participated in providing preference assessments over our dataset, in the form of an annotation survey (Appendix 11) where they were shown 15 prompts from the dataset, each of which included 1 response for each of the 5 strategies. Participants rated both the output and the strategy associated with the output in terms of appropriateness (on a scale 1 - 5). Combined with 5 attention checks, participants completed a total of 80 sub-tasks with a median time of 30 minutes. For the annotation portion, we recruited a total of 544 participants to cover the annotation on train and dev sets across two model types, guaranteeing 2 annotations per group per scenario.
For the end-to-end evaluation, we recruited additional annotators from each group, who assessed outputs produced after ICL alignment using a similar rating survey interface to the annotation process, instead framed as rating outputs from different systems rather than different strategies. For each end-to-end evaluation, in addition to participants who evaluated outputs aligned to their own group, we also set aside 1/3 of the users from each group to evaluate the outputs of alignment against all 4 groups aggregated as one large group to simulate a more collective value alignment setting (Section 4.7). We recruited a total of 240 participants to conduct the evaluation of prompting strategies (Section 4.5) and 120 participants for the end-to-end evaluation on the held-out test set (Section 4.6). Tasks were calibrated on Prolific to pay at a rate of \$12 USD/hour with the study design was reviewed and determined exempt by our IRB.
@@ 4.4 Results: Evaluating Retrieved Scenarios
Figure 3: Average error over preferences (measured by RMSE) comparing retrieved scenarios and observed ground preferences on the dev set. baseline uses similarity-only retrieval. spica uses the weighted group-informed metrics.
For our first evaluation, we examine whether group-informed metrics can improve the scenarios retrieved. To evaluate this, we look to the average error over preferences (RMSE) metric introduced in 3.1: A scenario with a lower error means that, for each response type in the scenario, preferences about its appropriateness are more similar to the actual observed preferences on that type of response for the input query. This evaluation is possible as we use queries from scenarios in the dev set, where preferences have also been collected over each response type. We follow the design of Spica to retrieve scenarios from the scenario bank (train) as ICL input examples. For the baseline condition, we only use the similarity metric over the prompts, while with Spica we utilize the additional group-informed metrics over preferences.
As indicated in Figure 3, we see that for both models, Spica is able to retrieve scenarios that had preferences more accurately matching those observed directly on the dev set. We also examined this effect at a per-group level, and find that while error is lowered across all groups, the magnitude of such change can differ (details of this we examine in Appendix A.1).
@@ 4.5 Results: Evaluating In-Context Prompting Strategies
In order to examine the effectiveness of ICL prompting setups (Section 3.3), we used human participants to evaluate the outputs produced models given each type of prompt while using the same Spica retrieval setup. Participants evaluated the outputs using an interface similar to that used during preference collection for scenario banks, however, instead of rating response strategies, participants rated 5 hypothetical AI systems (“System A - E”), each representing one configuration with a final control output produced by the model with no ICL alignment. Results are evaluated within-subjects, computing the difference between the rating of each condition’s aligned output and the control output, presented as “rating delta”.
As shown in Figure 4, we find that, for the gpt-4o model in a pluralistic alignment setting, the combination of contrastive response examples (C-R) proved to be the most effective (significant $p=0.030<0.05$ via ANOVA), on average rating $0.10$ points higher than the control across all groups. Additionally, given that scenario bank preferences may be collected only over high-level strategies, we also find positive instructions (P-I) to be the most effective prompt setup in this situation, producing ratings that average $0.07$ points higher (though not statistically significant). Given the results from this evaluation, the most effective prompting setup as observed on gpt-4o are P-I) and C-R. We use these two configurations in our final end-to-end evaluation on the test set, referring to the Spica with these prompt setups as Spica-I and Spica-R respectively.
We also examined these ICL prompts on the llama3-8b model. However, here none of the setups provided reliable alignment outcomes. In the pluralistic alignment task, we found no statistically significant differences between outputs, and average change in ratings were small or negative. We hypothesize that the smaller size of the llama3-8b model we used may be less capable at generalizing from ICL-style alignment.
@@ 4.6 Results: Evaluating End-to-End Alignment Outputs
Given the above findings on configurations, we conducted end-to-end evaluation that generate outputs for a held out test set of user queries. For comparison, we test a baseline ICA setup, where retrieval only considers semantic similarity with ICL prompt incorporating the highest rated response in each scenario as the demonstration example. For Spica, we use the two prompting strategies from above: Spica-I and Spica-R.
In Figure 5, we present results for the final end-to-end evaluation. We observe that on llama3-8b, none of the ICL alignment approaches performed well compared to the unaligned control, with the highest Spica-R averaging +0.005 / 5 rating points above control and no statistically significant differences between conditions overall or at a group-level. On gpt-4o, ICL alignment is effective, with the best system Spica-R performing better +0.072 / 5 points better than the unaligned control.
While considering all groups collectively, no condition was significantly better, if we look at each group individually, we find that for the rep-nrel group, Spica-R did result in statistically significant +0.16 points higher performance compared to baseline (within subjects paired t-test, $p=0.044<0.05$), with the rep-rel group also seeing a borderline significant improvement (within subjects paired t-test, $p=0.051$) of +0.16 points. Considering that recent work Rozado (2024) has found many current LLMs to favor liberal values, this result suggests that pluralistic alignment via Spica provides benefits primarily by improving alignment for traditionally disadvantaged groups.
Further examining alignment outcomes at a group level (Figure 9) we also find support that Spica can lead to more equitable outcomes across groups: With baseline on gpt-4o, we find that while the dem-rel and dem-nrel groups prefer aligned outputs (seen as +0.11, and +0.13 points over control), the rep-rel and rep-nrel groups end up preferring the original outputs (observed as -0.07, and -0.11 rating points compared to control). This discrepancy between groups is statistically significant for the minority group of non-religious Republican identifying (rep-nrel) participants (unpaired t-test between groups, $p=0.031$ and $p=0.049$). However, with Spica-R, all groups now prefer aligned aligned outputs (+0.10, +0.05, +0.09, +0.05) and we no longer see any statistically significant difference between groups in terms of this preference. Considering that in all conditions, ICL examples draw from each group’s own preference annotations (values), this supports the importance of considering group characteristics (norms) as a part of the retrieval process.
@@ 4.7 Results: Comparing Pluralistic versus Collective Alignment
One final aspect we wanted to explore at the end, was whether similar characteristics from pluralistic alignment carry over to a traditional collective value alignment setting. To accomplish this, in the evaluations for Section 4.4 and Section 4.6, we also included an additional type of system output—outputs aligned against all—where we used the same setup for each condition, with the only difference being that instead of having group-specific preferences, we define an “all” group that aggregates annotated preference judgments from all 4 groups. During evaluation, a portion of participants (Section 4.3) from each group would then be given this collectively aligned output to rate instead of group-specific alignment outputs.
In Figure 6, we see that unlike the pluralistic setting, when aligning to collective values of this super-group, optimal ICL prompt strategies on gpt-4o now significantly favor instructions (P-I and C-I) over example responses (llama3-8b still cannot be effectively aligned with either).
On the end-to-end evaluation of test set queries, while Spica-R is no longer an optimal prompt style for gpt-4o, we do see that group-informed retrieval metrics from Spica also no longer seem to provide any benefit, even slightly under-performing baseline retrieval. This is unsurprising though, as we would not expect pluralistic group-informed metrics to contribute significantly to the selection of scenarios for a “group” with few shared norms. Instead, now we find that the ICL prompt style becomes the most important factor, with gpt-4o favoring instructions and llama3-8b favoring example responses (used by both baseline and Spica-R), with this trend largely carrying over even across the participant’s own group affinity differences. This suggests that for traditional collective alignment, baseline retrieval setups are sufficient.
5 Discussion
5.1 The Role of Pluralistic Versus Collective Value Alignment
Considering the results of our experiments, there is a discussion to be had around the role of pluralistic versus collective value alignment. From prior work, we have seen how existing models can favor the values and norms of their designers and of majority populations Buyl et al. (2024); Rozado (2024), and how traditional alignment setups can be insufficient when addressing the challenges introduced by pluralistic alignment. However, while the groups in our evaluation were selected to illustrate how norms and values can differ, it is not hard to imagine how pushing for pluralistic alignment along demographic and ideological lines could lead to seemingly happier groups and communities at the cost of increased social fragmentation. Indeed, processes for consensus during collective value alignment Bai et al. (2022) can be means for society to find or build common values across groups.
Ultimately, we see a role for both pluralistic and collective value alignment in the broader landscape for building better socially-situated LLMs. Building from measurements around social norms like those used in Spica, there are opportunities for identifying pluralistic groups that reflect healthy diversity within society rather than simply being divisive. At the same time, collective value alignment processes, especially ones involving democratic inputs, remain important tools to bridge divides and find common ground.
5.2 Towards More Efficient Ways to Map of Group Values and Norms
In this work, we constructed scenario bank scenarios by drawing prompt data from an existing alignment dataset. While this can be a simple way to bootstrap the Spica framework, it is also not particularly efficient when it comes to mapping out meaningful norms and values of a group. In practice, we found that scenarios created in this way often end up illustrating similar values, leading to duplicated efforts during the annotation process for limited benefit. One avenue for improving the efficiency around eliciting group values that has been proposed by prior work is to engage group members in actively probing their own values through human-in-the-loop interactive systems Klingefjord et al. (2024). While current systems have primarily focused on more abstract value statements, similar designs could be adopted for groups to interactively expand upon scenario banks guided by their own exploration of the space of scenarios relevant to them. We believe there are many opportunities here for both new interactive system designs around collaboratively synthesizing scenarios, as well as new metrics that dynamically surface regions of scenarios that have been overlooked by each group.
5.3 Generalizing to Non-Discrete Preferences
In the specific implementation presented in this work, we apply Spica primarily in a discrete setting, defining preference functions through rating scores and responses through either concrete text or high-level strategies. While we discuss these limitations, we do also note that in practice alignment often applies to continuous settings: prompts, responses, and preferences all lie in a continuum.
While we do not explore this, the ideas introduced in Spica can largely generalize to non-discrete settings with few modifications. For example, while prompts and responses are currently sampled discretely, a non-discrete version of Spica may make use of embeddings and diffusion-based generative models Li et al. (2022) to examine regions of a continuous generation space. The measures in Spica could also be applied in settings where continuous preferences can be acquired, with concepts like “stability” and “contrast” potentially being captured through distribution-level properties like divergence and kurtosis.

### v2

@@ 3.2.3 Learning Metric Weights
While our metrics encode salience of scenarios for a specific group, we still need to balance this with the general relevance of scenarios to the input. In Spica, we do this by taking a linear weighted combination of the introduced metrics and a traditional similarity score (distance): $\bar{d}(x,x^{\prime})=w_{d}\cdot d(x,x^{\prime})+w_{s}\cdot g_{\text{stability}}(x^{\prime})+w_{c}\cdot g_{\text{contrast}}(x^{\prime})+c$.
As optimal weighting is likely to vary across groups, we empirically find these weights. Looking to Section 3.1.1, we note that the desirability of $x^{\prime}$ as an example given input $x$ can be assessed by the expected preference mismatch $\mathbb{E}_{y,y^{\prime}}[L(r(x,y),r(x^{\prime},y^{\prime}))]$. Thus for the final metric, we can compute this loss and minimize using linear regression $\bar{d}(x,x^{\prime})=\mathbb{E}_{y,y^{\prime}}[L(r(x,y),r(x^{\prime},y^{\prime}))]$. We note that the above equation considers only the best ($k=1$) example, with larger sets of ${x^{\prime}}$ possible by modifying the expression to include the loss for each additional example.
@@ 3.3 In-Context Learning Prompts for Retrieved Scenarios
Because retrieved scenarios contain preference distributions across multiple responses (or strategies), different setups for integrating scenarios as demonstrations are likely to produce different model outputs. ICL prompt designs have been extensively studied by prior works Sun et al. (2023); Higginbotham and Matthews (2024); Hao et al. (2022), so in this work we primarily explore new configurations enabled by the scenario bank. For one, preference distributions from scenario banks allow ICL examples to include multiple responses to illustrate more of the preference distribution: Rather than traditional retrieval which selects a Positive example of a good response, in Spica, we can select Contrasting examples that include both illustrations of a most preferred response as well as one that is least preferred. Additionally, the organization of responses into response classes means that scenario banks can provide either concrete examples of Response text, or higher level Instructions that lead to producing a response in that response class. Altogether, this creates 4 combinations of prompt setups that we can use: P-I, C-I, P-R, and C-R. We discuss our implementation and evaluation in the sections that follow.
@@ 4 Experiments and Results
To evaluate Spica, we set up a pluralistic alignment task involving 4 demographically constructed groups, and assess how well a Spica workflow is able to align model outputs to preferences of each group compared to a baseline approach that only considers semantic similarity.
@@ 4.1 Dataset and Scenario Bank Construction
For our evaluation alignment task, we constructed a set of queries (which define the topics to provide alignment on) by drawing from an existing set of challenging alignment situations based on prompts observed in conversations on the PRISM dataset Kirk et al. (2024). PRISM engaged human participants to interact with LLMs by naturally starting conversations with 3 types of guidance meant to invoke conversations around more challenging and complex topics: “unguided”, “values guided”, or “controversy guided”. We observed that of the 3 types of guidance, unguided conversations primarily resulted in simple informational requests which are not particularly controversial in the context of pluralistic alignment, so we opted to drop conversations of this type. Among the remaining conversations, we randomly selected a subset, split into 3 slices: retrieval (train, $n=360$), weight optimization and selection of ICL prompt setups (dev, $n=150$), and evaluation hold-out (test, $n=75$).
As PRISM responses are created by existing collective-value-aligned models, they do not cover desirable behaviors for all groups. Instead, we follow Section 3.1 and construct new responses ourselves based on several classes of common model behaviors (Appendix A.7.1). To capture the stochastic nature of model outputs, we generate 3 responses in each class.
@@ 4.2 Models and Similarity Metric
For our experiments, we tested the quality of retrieval-based ICL alignment using one open-source (llama3-8b) and one closed-source model (gpt-4o-2024-05-13) as the base model. llama3-8b [^2] We considered using 70b, but could not reliably run inference due to memory limitations of available hardware. inference was conducted using a locally hosted instance of Ollama [^3] https://ollama.com/. With both models, we applied the same prompts to generate responses attached to scenario bank queries and to conduct in-context alignment (Appendix A.7.2). As our goal is to evaluate the additional metrics we introduced, we kept the semantic similarity measurements constant across all models and conditions, using values derived by computing the cosine similarity between embeddings generated by text-embedding-3-large from OpenAI.
@@ 4.3 Pluralistic Groups and Human Annotation Setup
We define four groups in the form of demographic slices drawn from the US population: partisan political affiliation (“republican” or “democrat”), and self-reported regular participation in religious activities (“yes”—rel or “no”—nrel). Our choice of these features is based on similar factors that were salient for opinions around AI Zhang and Dafoe (2019) along with practical considerations around demographic splits that we could reliably recruit on our crowd work platform, Prolific.
Annotators in each group participated in providing preference assessments over our dataset, in the form of an annotation survey (Appendix 11) where they were shown 15 prompts from the dataset, each of which included 1 response for each of the 5 model behavior classes. Participants rated both the output and the description of the behavior class associated with the output in terms of appropriateness (from 1–“inapproprate” to 5–“appropriate”). Combined with 5 attention checks, participants completed a total of 80 sub-tasks with a median time of 30 minutes. For the annotation portion, we recruited a total of 544 participants to cover the annotation on train and dev sets across two model types, guaranteeing 2 annotations per group per scenario. In the end-to-end evaluation (Section 4.6), we recruited separate annotators from each group, who assessed outputs produced after ICL alignment. Annotators used the same survey interface, though they rated outputs produced by different conditions rather than outputs by response class. For each end-to-end evaluation, we set aside 1/3 of the users from each participant group to evaluate the outputs of collective alignment (Section 4.7) which uses aggregated rather than group-specific preferences. We recruited a total of 240 participants to conduct the evaluation of prompting strategies (Section 4.5) and 120 participants for the end-to-end evaluation on the held-out test set (Section 4.6). Tasks were paid at a rate of \$12 USD/hour, and the study design was deemed exempt by our IRB.
@@ 4.4 Results: Evaluating Retrieved Scenarios
Figure 3: Average error over preferences (measured by RMSE) comparing retrieved scenarios and ground truth on the dev set. baseline uses similarity-only retrieval. spica uses weighted group-informed metrics.
For our first evaluation, we examine whether group-informed metrics result in the retrieval of better examples. In 3.1 we noted that, for a new user query, retrieving a scenario whose known behavior preference distributions better matched the post-hoc observed behavior preferences of responses to the query would indicate a desirable outcome. We measure this mismatch (or error) following the approach outlined in Section 3.1.1. Since multiple participants provide behavior preferences $r_{p}$ (both in the scenario bank and as part of the ground truth on the dev set), we take the average across all pairwise error measurements between the two.
After tuning the weights for metrics as noted earlier in Section 3.2.3, we find that with both models, Spica retrieves scenarios that had preference distributions more accurately matching the observed ground truth distributions on the dev set (Figure 3). While this result should not be surprising, it does indicate that for pluralistic alignment, there was room for improvement on the retrieval metric. We also note that at a per-group level, while error is lowered across all groups, the magnitude of this difference varies between groups (Appendix A.1).
@@ 4.5 Results: Evaluating In-Context Prompting Strategies
In order to examine the effectiveness of ICL prompting setups (Section 3.3), we used human participants to evaluate the outputs produced by models given each type of prompt while using the same Spica retrieval setup. Participants evaluated the outputs using an interface similar to that used during preference collection for scenario banks. However, instead of rating response strategies, participants rated on a 1 - 5 scale 5 hypothetical AI systems (“System A - E”), each representing one configuration with a final control output produced by the model with no ICL alignment. As we used a within-subjects design, we measured alignment outcomes by computing the difference between each participant’s rating of an aligned output (each condition) and the reference control output, which we report as the “rating delta”.
We find (Figure 4) that for the gpt-4o model in a pluralistic alignment setting, the combination of contrastive response examples (C-R) proved to be the most effective (significant $p=0.030<0.05$ via ANOVA), on average rating $0.10$ points higher than the control across all groups. We also found that positive instructions (P-I) were also somewhat (though not significantly) more effective, resulting in $0.07$ point higher ratings. Using the same prompts with the llama3-8b model, we did not find any setup that provided reliable improvements to model outputs, with no significant differences observed between conditions and differences small or negative. We hypothesize the smaller llama3-8b model may have contributed to less capability when generalizing via ICL-style alignment.
Overall, we found that P-I and C-R were most promising, and we used these two configurations in our end-to-end evaluation on the test set. We will refer to these as Spica-I and Spica-R respectively.
@@ 4.6 Results: Evaluating End-to-End Alignment Outputs
We conducted an end-to-end evaluation that generates outputs for a held out test set of user queries. As a baseline, we used a traditional ICA setup where retrieval only uses semantic similarity, and the ICL prompt only incorporates the highest rated response for each scenario retrieved. For Spica, we use the two best prompt setups from Section 4.5, Spica-I and Spica-R. As seen in Figure 5, we find that for gpt-4o, ICA was generally effective, with Spica-R being the best system, performing +0.072 / 5 points better than the control, while on llama3-8b, ICL alignment produced marginal results, with Spica-R still being the best system but only averaging +0.005 points above baseline.
When considering all groups, no condition was significantly better. However, if we look at each group (Section A.4), we find that for the rep-nrel (Republican, non-religious identifying) group, Spica-R resulted in a statistically significant +0.16 points higher performance compared to baseline (within subjects paired t-test, $p=0.044<0.05$), with the rep-rel group also seeing an improvement (within subjects paired t-test, $p=0.051$) of +0.16 points. Given recent work Rozado (2024) finding many LLMs favor liberal values, this result suggests that pluralistic alignment via Spica benefitted alignment primarily by improving outcomes for traditionally disadvantaged groups.
Further examining alignment at a group level, we also find support that Spica can lead to more equitable outcomes across groups (Figure 8); with baseline on gpt-4o, we find that while the dem-rel and dem-nrel groups prefer our aligned outputs (seen as +0.11, and +0.13 points over control), the rep-rel and rep-nrel groups end up preferring the original outputs (observed as -0.07, and -0.11 rating points under control). This discrepancy between groups is statistically significant for the minority group of rep-nrel participants (unpaired t-test between groups, $p=0.031$ and $p=0.049$). However, with Spica-R, all groups now prefer aligned outputs (+0.10, +0.05, +0.09, +0.05) and we no longer see any statistically significant difference between groups in terms of this preference. Despite ICL examples themselves drawing from each group’s own preferences in all conditions, this result indicates that retrieving the right examples (by considering group norms) can improve equitable outcomes across groups.
@@ 4.7 Results: Comparing Pluralistic versus Collective Alignment
If retrieval metrics based on group norms were helpful for alignment, why have more traditional collective alignment processes not used them? To investigate this, we combined all 4 groups into one collective group and provided an additional output (all) during the evaluations for Section 4.4 and Section 4.6 produced by applying Spica on these collective preferences. Unsurprisingly, we found (Section A.5) that Spica’s metrics contributed little in this collective alignment setting, with traditional similarity-based retrieval being largely sufficient, suggesting a reason why group-informed metrics may not have been explored by past works.
5 Conclusions and Discussion
In this work, we propose Spica as a new framework to support pluralistic alignment. Through evaluations, we find that group-informed metrics coupled with the scenario bank and ICL prompts in Spica contributed to improving pluralistic alignment, primarily by supporting groups that are traditionally disadvantaged.
Pluralistic Versus Collective Values
From prior work, we have seen how existing models can favor the values and norms of their designers and of majority populations Buyl et al. (2024); Rozado (2024) in collective alignment settings. With our work on Spica, we also present a path towards supporting pluralistic alignment towards individual groups. However, focusing on pluralistic alignment alone can lead to divides along demographic and ideological lines, furthering social fragmentation. Ultimately, we believe there should be a balance between striving for common ground through collective alignment Bai et al. (2022), and accommodating diverse views through pluralistic alignment.
Efficiently Mapping Group Values and Norms
In this work, we built our scenarios by drawing from existing conversation data. However, this is not a very efficient way to map group values—many user queries may not have controversial model behaviors and even controversial conversations end up covering similar points of contention. With the increased capability of models, we believe future work may be able to dynamically elicit group values much more efficiently through interactive LLM-backed agents engaging with groups in human-in-the-loop refinement and synthesis processes Klingefjord et al. (2024) that could produce scenarios that are either better demonstrations of values and norms or more controversial to ground ambiguous decision bounds.